OpenStudy (anonymous):

What is an equation of the line in slope-intercept form? m = 2/7 and the y-intercept is (0, –12)

OpenStudy (anonymous):

The slope-intercept form is generally written as y=mx+c where c is the y-intercept ie the point where the line intersects the y-axis. Substituting the values in the equation : y = (2/7)*x + (-12) = (2/7)x -12

OpenStudy (lstrasmann):

\[y = \frac{ 2 }{ 7 }x - 12\]